OpenStudy (anonymous):

whats the integral of (lnt)^2

OpenStudy (watchmath):

Integration by parts \(u=\ln t\) then \(du=2\ln t /t\) \(dv=dt\) then \(v=t\) It follows that the integral is equal to \(t\ln t-\int 2\ln t\,du\) Then use integration by parts one more :D